A series RLC circuit is being driven by a sinusiod
Answer #1

IRES = V/R = 1.25A

I0 = IRES/2 = 0.625 A

we know that

I = V/Z

therefore

Z = 2R

XL-Xc = R sqrt 3

XL--Xc = 207.84 ohm

wL - 1/wC = 207.84

w^2(LC) -0.00311769 w -1

w = 2360 rad/sec